The sum of three consecutive numbers is 27 and the sum of their square is 293. Find the numbers.
Answer by LinnW(1048)

x the first no
x + 1 the second number
x + 3 the third number
(x)+(x+1)+(x+2)=27
Let's see if there is a unique solution.
3x + 3 = 27
subtract 3 from each side
3x = 24
x = 8
so the numbers are 8,9,10
It could be that the sum of their squares is 293
64 + 91 + 100 ?= 293
For this problem there appears to be no solution
since we cannot satisfy both requirements.